# Molecular Communication Channel as a Physical Reservoir Computer

### v1

@@ Abstract
Molecular Communication (MC) channels inherently possess significant memory and nonlinear dynamics due to diffusion and receptor kinetics, often posing challenges for reliable data transmission. This work reconceptualizes these intrinsic properties as computational resources by framing a canonical point-to-point MC channel, consisting of ligand diffusion and reversible ligand-receptor binding at a spherical receiver, as a physical reservoir computer (PRC). We utilize the time-varying fraction of bound receptors as the reservoir’s internal state, employing time-multiplexing to generate high-dimensional virtual nodes without explicit recurrence. Only a linear readout layer is trained via ridge regression. Through deterministic mean-field modeling and particle-based spatial stochastic simulations, we demonstrate the MC system’s capability for complex temporal processing by successfully performing next-step prediction on standard chaotic time-series benchmarks (Mackey-Glass and NARMA10). Performance, quantified by Normalized Root Mean Square Error (NRMSE), exhibits a non-monotonic dependence on key system parameters (receptor kinetic rates, diffusion coefficient, transmitter-receiver distance), revealing optimal operational regimes. These findings validate the potential of using MC channel as effective and low-complexity computational substrate.
Index Terms:
Reservoir Computing, Molecular Communications, Ligand-Receptor Interactions
@@ I Introduction
Reservoir Computing (RC), introduced independently as the echo‑state network (ESN) by Jaeger [1] and the liquid‑state machine (LSM) by Maass et al. [2], is a class of recurrent neural network models in which an untrained, nonlinear dynamical system (the reservoir) projects an input into a high‑dimensional state space, while only a linear readout is trained, typically via ridge regression. This separation of representation and learning obviates back‑propagation through time, providing fast, data‑efficient training while retaining the rich temporal processing power of recurrent dynamics. As a result, RC has proven highly effective for time-series prediction tasks [3].
@@
Physical Reservoir Computing (PRC) generalizes this principle by exploiting the native dynamics of physical substrates, e.g., photonic [4], electronic [5], spintronic [6], and mechanical systems [7], as reservoirs, often in delay‑based architectures that implements large virtual networks through time-multiplexing of a single nonlinear node [8, 9]. Such implementations greatly reduce hardware complexity while achieving performance on par with digital RC, and have catalyzed interest in neuromorphic edge computing [10] and embodied intelligence [11].
@@
Recent work has begun to evaluate biological systems as reservoirs, capitalizing on their inherently rich and adaptive dynamics. Pioneering experiments with neuron cultures [12], brain organoids [13], bacteria [14], and plants [15] demonstrate that living systems can be co-opted or integrated into artificial computing frameworks to execute externally defined computational tasks. Likewise, chemical reaction networks, which underlie many fundamental biochemical processes, have been investigated for RC, demonstrating the potential to exploit complex molecular interactions for sophisticated information processing beyond their native biological context [16].
@@
Central to the exceptional information processing capabilities of biological systems is biochemical signaling, which enables coordination and adaptation across multiple scales, critical for their survival [17]. Building on these naturally evolved mechanisms, researchers have long studied engineered molecular communication (MC) systems as alternative to conventional electromagnetic (EM) communications to enable networks of micro/nanoscale devices via the exchange of precisely controlled chemical signals [18]. This line of work has fostered the vision of the Internet of Bio-Nano Things (IoBNT), in which biological entities and artificial micro/nanodevices interface seamlessly for next-generation biomedical applications [19]. Extensive theoretical research and recent experimental work in this field have particularly focused on overcoming the unique challenges of MC channels that are different than EM communications, including significant channel memory effects (inter-symbol interference) and nonlinear receiver responses (involving receptor saturation) in addition to biochemical noise [20, 21]. These complexities turned out to be particularly demanding for small, resource-limited devices, yet they also hint at untapped computational potential that biological systems may be inherently exploiting.
@@
In this paper, we reconceptualize MC channel as a physical reservoir computer, revealing its potential for information processing and implementing machine learning tasks. Specifically, we study a simple MC channel model, which is adequately representative of both biological and synthetic MC systems, consisting of a point transmitter and a spherical receiver whose surface is functionalized with ligand receptors. By adopting a time-multiplexed RC framework, we leverage the inherent channel memory arising from diffusion and the nonlinear ligand-receptor interactions at the receiver, both of which are critical for computational performance. Through numerical analysis and particle-based spatial stochastic simulations, we demonstrate that MC channels can successfully perform chaotic time-series prediction tasks with minimal training, as benchmarked by standard RC evaluations on NARMA10 and Mackey-Glass time-series. We systematically investigate how various MC system parameters, including the number of transmitted molecules, transmitter-receiver distance, diffusion coefficient, and ligand-receptor kinetic rates, affect computing performance.
@@
Our results provide evidence that even simple MC channels can be harnessed for non-trivial time-series tasks by virtue of their intrinsic diffusion-driven memory and receptor-mediated nonlinearities. Beyond contributing to the fundamental understanding of information-processing capabilities of biological systems at the elemental level of their communication links, this computing perspective on MC channels can open new avenues for bio-inspired and bio-integrated computing architectures. Potential applications range from early disease diagnosis analyzing time-varying biomarker profiles to channel equalization in artificial MC networks and IoBNT [22]. By demonstrating how the channel’s inherent dynamics can be systematically exploited for information processing, this study lays the groundwork for MC systems to serve not only as physical channels for communication but also as powerful computational substrates in their own right.
@@
The paper is organized as follows. Section II reviews molecular communication and physical reservoir computing. Section III formalizes the molecular‑communication channel as a physical reservoir computer and describes the modeling and training methodology. Section IV presents deterministic and stochastic simulation results with parameter studies. Section V concludes and outlines future directions.
@@ II-A Molecular Communications
Bio-inspired MC emerged as a promising approach to realize networks of micro/nanoscale devices, such as biosensors and microrobots, and to seamlessly interface these engineered networks with biological systems. Such integration underpins the vision of the IoBNT, where MC mediates collaborative tasks such as targeted drug delivery and continuous health monitoring [23]. By exploiting mechanisms similar to cellular signaling, MC provides bio-compatible, low-energy connectivity in environments where conventional EM methods become impractical [24].
From early theoretical studies on capacity and error rates, the MC field has advanced to increasingly sophisticated experiments [21, 25, 26, 27]. Researchers have identified unique challenges that distinguish MC from EM communications, including severe inter-symbol interference (ISI) due to slow diffusion, and nonlinearity plus memory effects from stochastic receptor binding and other reaction kinetics. Practical MC transceivers at the microscale also face tight constraints in energy, size, and computational resources, motivating new strategies in encoding, detection, and synchronization [20]. To address these issues, data-driven and machine learning methods have emerged, with neural networks learning the input-output mapping of MC channels directly from data and thus coping better with time-varying conditions and nonlinear dynamics [28].
In parallel, an expanding line of research investigates MC for in situ artificial intelligence (AI) with living cells and engineered biomolecular processes, commonly referred to as wetware AI. Two complementary strategies have emerged: top-down approaches that leverage existing cellular functionalities, and bottom-up approaches that build computing frameworks from fundamental chemical reactions. In the former category, wet-neuromorphic computing was introduced by repurposing gene regulatory networks (GRNs) for pattern recognition [29], while Wet TinyML concept applied GRN-based methods for energy-efficient neural-like computations within living cells [30]. Another study integrated biologically plausible spiking neurons into reservoir computing, using MC-based connectivity for distributed processing among cells [31]. A unifying framework for Molecular Machine Learning (MML) positioned microbial consortia in addition to GNRs within individual cells as native substrates for AI, with MC serving as the core mechanism for data exchange and biochemical learning [32].
@@ II-B Physical Reservoir Computing
In a standard RC scheme, the input vector $\mathbf{u}(t)$ drives the reservoir state $\mathbf{x}(t)\in\mathbb{R}^{N}$ through
@@
where $\mathbf{A}\in\mathbb{R}^{N\times N}$ is a fixed (typically random) recurrent weight matrix, $\mathbf{W}_{\mathrm{in}}$ projects the input into the reservoir, and $f(\cdot)$ is an element‑wise nonlinearity (Fig. 1(a)) [36]. The readout is obtained by a linear map
$\mathbf{y}(t)=\mathbf{W}_{\mathrm{out}}\mathbf{x}(t),$ (2)
with $\mathbf{W}_{\mathrm{out}}$ trained, often by ridge regression, to minimize the discrepancy between $\mathbf{y}(t)$ and a target signal. Because the internal weights $\mathbf{A}$ remain fixed, RC sidesteps the heavy cost of back‑propagating gradients through deep recurrent structures, a feature that makes it well‑suited to hardware realizations.
PRC replaces the software‑defined reservoir with the intrinsic dynamics of a physical substrate. Phenomena such as optical feedback, spin‑wave interactions, analogue electronic circuits, and even the interference patterns of water waves in a bucket implement the operations $\mathbf{A}\mathbf{x}(t)$ and $f(\cdot)$ in real time, while external sensors capture the instantaneous state $\mathbf{x}(t)$ (Fig. 1(b)) [9, 35]. Leveraging the medium’s native parallelism, bandwidth, and energy efficiency, PRC offers a neuromorphic alternative to digital simulations.
Time‑multiplexed reservoirs. High‑dimensional states can be generated with minimal hardware by time‑multiplexing a single nonlinear element (Fig. 1(c)) [8]. During each clock period of length $T$, the input segment is multiplied by a masking sequence $\mathbf{m}$, and the element’s output is sampled at $M$ equidistant instants. The resulting vector
@@
acts as an $M$‑node virtual reservoir whose evolution from one period to the next is governed by the masked input and the device’s nonlinear response [37].
A widely adopted form of time‑multiplexing employs an external feedback loop of delay $\tau$ around a single nonlinear node (e.g., a semiconductor laser or an analogue circuit). Partitioning the delay into $M$ sub‑intervals of length $\Delta=\tau/M$ yields $M$ virtual nodes [9, 38]. Sampling the output every $\Delta$ seconds constructs $\mathbf{x}(t)$, while the recirculating signal supplies fading memory that underpins tasks such as speech recognition and chaotic time‑series forecasting. Additional delay lines or delayed taps at the input or readout can further enrich the dynamics or extend the memory horizon without increasing the number of physical components [39].
@@ III Molecular Communication Channel as a Physical Reservoir
We consider a single-link MC system consisting of a point transmitter and a spherical receiver, as illustrated in Fig. 1(d). The transmitter release a prescribed number of ligands (molecules) instantaneously at prearranged times into a fluidic medium, where they propagate by free diffusion. A fraction of these ligands eventually arrives near the receiver and can bind reversibly to receptors on its surface under given reaction kinetics. Below, we describe how this widely studied MC system model can be reframed as a physical reservoir computer by regarding the time-varying fraction of bound receptors as the high-dimensional dynamical state.
Let the transmitter release $I(n)$ molecules at each symbol time $nT$, where $n\in\{0,1,2,\dots\}$, and $T$ is the symbol interval. For simplicity, we assume the receiver is at distance $x_{R}\gg r_{\mathrm{rx}}$ from the transmitter at the origin, such that the receiver’s finite radius $r_{\mathrm{rx}}$ can be neglected in the channel impulse response. Under three-dimensional (3D) free diffusion with diffusion coefficient $D$, the ligand concentration $c(\mathbf{x},t)$ obeys Fick’s second law:
@@
with the far-field boundary condition $c(\mathbf{x},t)\to 0$ as $\|\mathbf{x}\|\to\infty$. The impulse response at distance $x_{0}$ for a single pulse released at $t=0$ is the 3D diffusion kernel
$h(t)\;=\;\frac{1}{(4\pi Dt)^{3/2}}\exp\!\Bigl{(}-\frac{x_{0}^{2}}{4Dt}\Bigr{)}\quad(t>0).$ (4)
Hence, when the transmitter emits $I(n)$ molecules at times $nT$, the local ligand concentration at the receiver (treated as a point in the far-field approximation) is given by
$c_{T}(t)\;=\;\sum_{n=0}^{\infty}I(n)\,h\bigl{(}t-nT\bigr{)}\,u\bigl{(}t-nT\bigr{)},$ (5)
where $u(\cdot)$ is the unit step function enforcing causality. Since $h(t)$ has a infinitely long tail (though it decays rapidly), successive transmissions introduce ISI. For practical numerical simulations, this tail is often truncated to a finite channel memory length [40].
Figure 2: (a) An instance from particle-based stochastic simulations of the MC channel with Smoldyn. Receptor occupation ratios for numerical and stochastic simulations of the MC system for (b) Mackey-Glass series input, (c) NARMA10 Series input, (d) Mackey-Glass series input, post filtering (smoothing), (e) NARMA10 series input, post filtering.
The spherical receiver is equipped with $N_{R}$ identical, monovalent ligand receptors. Each receptor can be either unbound or bound to one ligand, with reversible reaction:
@@
where $L$ is the free ligand, $R$ is an unbound receptor, and $LR$ is a ligand-receptor complex, i.e., bound receptor. At the mean-field level, and assuming the local free ligand concentration is not significantly depleted by bound receptors, the time evolution of the mean fraction of bound receptors, $b(t)=n_{b}(t)/N_{R}$, can be approximated by
$\frac{d\,b(t)}{dt}\;=\;k_{\mathrm{on}}\,c_{T}(t)\,\bigl{[}1-b(t)\bigr{]}\;-\;k_{\mathrm{off}}\,b(t),$ (6)
where $k_{\mathrm{on}}$ and $k_{\mathrm{off}}$ are the forward (binding) and reverse (unbinding) rate constants, respectively, and $c_{T}(t)$ is the local ligand concentration at the receiver.
Although the exact time-varying variance of $b(t)$ under such nonstationary input is typically not available in closed form (due to correlated ligand arrivals and receptor occupancy fluctuations), we use particle-based spatial stochastic simulations (with Smoldyn (Fig. 2) [41]) in Section IV to account for this stochasticity. For the present analytical treatment, (6) provides a deterministic approximation to the mean behavior of the ligand-receptor interactions.
[…]
In RC, one aims to transform an input time series $\{u(n)\}$ into a high-dimensional dynamical state $\mathbf{x}_{n}$ through a fixed nonlinear system (the reservoir), and then train only a linear readout for the target computational task. Here, we map the input $u(n)$ to the number of released molecules $I(n)$ at each symbol time $nT$:
@@
where $\alpha$ is a scaling factor. The channel then imposes a complex, memory-rich temporal response on these inputs (due to both diffusion and receptor binding).
To extract a sufficiently high-dimensional state from the scalar output $b(t)$, we perform time-multiplexing. That is, within each interval $[nT,\,(n+1)T]$, we sample $b(t)$ at $M$ distinct time offsets $\{\tau_{1},\tau_{2},\dots,\tau_{M}\}$:
@@
This sampling effectively creates $M$ virtual nodes per symbol. Unlike a typical optoelectronic delay-based reservoir, however, we do not require a separate external feedback loop; rather, the channel’s intrinsic ISI provides the dynamical memory. The “masking” function typically used in conventional time-multiplexed RC is replaced here by the MC channel impulse response, which shapes the input in time to eventually generate diverse node responses upon sampling.
We train a linear readout to map from the reservoir state $\mathbf{x}_{n}$ to the desired output $y_{n}$ (e.g., a time-series prediction). Let $\mathbf{W}_{\mathrm{out}}$ be the $M$-dimensional readout weight vector, then
@@
With $N_{\mathrm{train}}$ training samples $\{(\mathbf{x}_{n},\,y_{n})\}$, we solve:
@@
where $\mathbf{X}$ is the $N_{\mathrm{train}}\!\times\!M$ matrix whose $n$-th row is $\mathbf{x}_{n}^{\top}$, $\mathbf{y}$ is the vector of target outputs $y_{n}$, and $\lambda$ is a regularization parameter (ridge regression). The closed-form solution is then given by the Moore–Penrose pseudoinverse:
@@
Hence, only $\mathbf{W}_{\mathrm{out}}$ is optimized, and the underlying channel and binding dynamics remain unmodified.

### v2

@@ Abstract
Molecular Communication (MC) channels are characterized by significant memory and nonlinear dynamics arising from diffusion and receptor kinetics. While often viewed as impairments to reliable data transmission, this work introduces a paradigm shift by reconceptualizing these intrinsic physical properties as computational resources. We frame a canonical point-to-point MC channel, comprising ligand diffusion and reversible ligand-receptor binding at a spherical receiver, as a Physical Reservoir Computer (PRC). Utilizing deterministic mean-field modeling and particle-based spatial stochastic simulations, we demonstrate the MC system’s inherent capability for complex temporal information processing on standard chaotic time-series benchmarks. We comprehensively evaluate performance using both task-specific Normalized Root Mean Square Error (NRMSE) and the task-independent Information Processing Capacity (IPC). Our results reveal a non-monotonic dependence of computational power on key biophysical parameters (receptor kinetic rates, diffusion coefficient, and transmitter-receiver distance), identifying optimal operational regimes where memory and nonlinearity are balanced. These findings establish the MC channel as a viable computational substrate, paving the way for novel architectures in wetware artificial intelligence.
@@ I Introduction
Reservoir Computing (RC), introduced independently as the echo‑state network (ESN) by Jaeger [1] and the liquid‑state machine (LSM) by Maass et al. [2], is a class of recurrent neural network (RNN) models in which an untrained, nonlinear dynamical system (the reservoir) projects an input into a high‑dimensional state space, while only a linear readout is trained, typically via ridge regression. The separation of representation and learning in RC bypasses the need for back-propagation through time, enabling fast, data‑efficient training while retaining the computational power of recurrent dynamics. As a result, RC has proven highly effective for challenging time-series prediction tasks [3].
@@
Physical Reservoir Computing (PRC) generalizes this principle by exploiting the native dynamics of physical substrates, e.g., photonic [4], electronic [5], spintronic [6], and mechanical systems [7], as reservoirs, often in delay‑based architectures that implement large virtual networks through time-multiplexing of a single nonlinear node [8, 9]. Such implementations greatly reduce hardware complexity while achieving performance on par with digital RC, and have catalyzed interest in neuromorphic edge computing [10] and embodied intelligence [11].
@@
Recent work has evaluated biological systems as reservoirs, capitalizing on their inherently rich, adaptive, and energy-efficient dynamics. Pioneering experiments with neuron cultures [12], brain organoids [13], bacteria [14], and plants [15] demonstrate that living systems can be co-opted or integrated into artificial computing frameworks to execute externally defined computational tasks. Likewise, chemical reaction networks, which underlie many fundamental biochemical processes, have been investigated for RC, demonstrating the potential to exploit complex molecular interactions for sophisticated information processing beyond their native biological context [16].
@@
Central to the exceptional information processing capabilities of biological systems is biochemical signaling, which enables coordination and adaptation across multiple scales, critical for their survival [17]. Building on these naturally evolved mechanisms, researchers have long studied engineered molecular communication (MC) systems as alternative to conventional electromagnetic (EM) communications to enable networks of micro/nanoscale devices via the exchange of precisely controlled chemical signals [18]. This line of work has fostered the vision of the Internet of Bio-Nano Things (IoBNT), in which biological entities and artificial micro/nanodevices interface seamlessly for next-generation biomedical applications [19, 20]. Extensive theoretical research and recent experimental work in this field have particularly focused on overcoming the unique challenges of MC channels that are different than EM communications, including significant channel memory effects (inter-symbol interference) and nonlinear receiver responses (involving receptor saturation) in addition to biochemical noise [21, 22]. These complexities turned out to be particularly demanding for small, resource-limited devices, yet they also hint at untapped computational potential that biological systems may be inherently exploiting.
@@
In this paper, we reconceptualize MC channel as a physical reservoir computer, revealing its potential for information processing and implementing machine learning tasks. Specifically, we study a canonical MC channel model, which is representative of both biological and synthetic MC systems, consisting of a point transmitter and a spherical receiver whose surface is functionalized with ligand receptors. We demonstrate that the inherent physical dynamics of this communication system, namely the channel memory from ligand diffusion and the nonlinearity from receptor kinetics, can be directly leveraged for complex computation. This is achieved by adopting a time-multiplexed RC framework to translate these continuous physical states into a high-dimensional computational state space with minimal overhead.
To characterize the computational power of the MC reservoir, we conduct a comprehensive parametric study, quantifying performance through two complementary lenses. First, we investigate task-specific predictive accuracy on standard chaotic benchmarks (Mackey–Glass and NARMA10). Second, we evaluate its task-independent computational power by analyzing the resulting Information Processing Capacity (IPC). This dual analysis provides critical insights as to how performance is fundamentally governed by key biophysical parameters, such as ligand-receptor kinetic rates, diffusion coefficient, and transmission distance, generating the first guidelines for optimizing MC systems for computation.
@@
Our results demonstrate that even simple MC channels possess the intrinsic capacity for non-trivial time-series processing. Beyond contributing to the fundamental understanding of information-processing capabilities of biological systems at the elemental level of their communication links, this computing perspective on MC channels can open new avenues for bio-inspired and bio-integrated computing architectures. Potential applications range from real-time analysis of time-varying biomarker profiles for early disease diagnosis to channel equalization in artificial MC networks and IoBNT [23]. By demonstrating how the channel’s inherent dynamics can be exploited for information processing, this study lays the groundwork for MC channels to serve not only as physical medium for unconventional communications but also as powerful computational substrates in their own right.
@@
The paper is organized as follows. Section II reviews the fundamentals of MC and PRC. Section III formalizes the MC channel as a physical reservoir computer and describes the modeling and training methodology. Section IV presents the performance evaluation, using both task-specific NRMSE and task-independent IPC to reveal the optimal operational regimes governed by the MC system parameters. Section V concludes the study and outlines future directions.
@@ II-A Molecular Communications
Bio-inspired MC has emerged as a key enabling technology for realizing networks of micro/nanoscale devices, such as biosensors and microrobots, and seamlessly interfacing them with biological systems [18]. This integration underpins the IoBNT vision, where MC mediates collaborative tasks such as smart drug delivery and continuous health monitoring [24]. By exploiting mechanisms analogous to cellular signaling, MC provides bio-compatible, low-energy connectivity in environments where conventional EM methods are impractical [25].
From early theoretical studies on channel capacity and error rates, the MC field has advanced to increasingly sophisticated experiments [22, 26, 27, 28]. Researchers have identified unique challenges that distinguish MC from EM communications, including severe inter-symbol interference (ISI) due to slow diffusion, and nonlinearity plus memory effects from stochastic receptor binding and other reaction kinetics. Practical MC transceivers at the microscale also face tight constraints in energy, size, and computational resources, motivating new strategies in encoding, detection, and synchronization [21]. To address these issues, data-driven and machine learning methods have emerged, with neural networks learning the input-output mapping of MC channels directly from data and thus coping better with time-varying conditions and nonlinear dynamics [29, 30].
In parallel, an expanding line of research investigates MC for in situ artificial intelligence (AI) with living cells and engineered biomolecular processes, commonly referred to as wetware AI. Two complementary strategies have emerged: top-down approaches that leverage existing cellular functionalities, and bottom-up approaches that build computing frameworks from fundamental chemical reactions. In the former category, wet-neuromorphic computing was introduced by repurposing gene regulatory networks (GRNs) for pattern recognition [31], while Wet TinyML concept applied GRN-based methods for energy-efficient neural-like computations within living cells [32]. Another study integrated biologically plausible spiking neurons into reservoir computing, using MC-based connectivity for distributed processing among cells [33]. A unifying framework for Molecular Machine Learning (MML) positioned microbial consortia in addition to GRNs within individual cells as native substrates for AI, with MC serving as the core mechanism for data exchange and biochemical learning [34].
@@ II-B Physical Reservoir Computing
In a standard RC scheme (Fig. 1(a)), an input vector $\mathbf{u}(t)$ drives the reservoir state $\mathbf{x}(t)\in\mathbb{R}^{N}$ through
@@
where $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the fixed (typically random) recurrent weight matrix, $\mathbf{W}_{\mathrm{in}}$ is the input weight matrix that projects the input into the reservoir, and $f(\cdot)$ is an element‑wise nonlinear activation function [38]. The readout is generated by a linear transformation of the reservoir state:
$\mathbf{y}(t)=\mathbf{W}_{\mathrm{out}}\mathbf{x}(t).$ (2)
The readout weights $\mathbf{W}_{\mathrm{out}}$ are trained, typically via ridge regression, to minimize the error between $\mathbf{y}(t)$ and a target signal. Since $\mathbf{A}$ and $\mathbf{W}_{\mathrm{in}}$ remain fixed, RC sidesteps the computationally intensive gradient descent required for training traditional RNNs, making it highly suitable for hardware implementation.
PRC replaces the software‑defined reservoir with the intrinsic dynamics of a physical substrate. Phenomena such as optical feedback, spin‑wave interactions, analogue electronic circuits, and even the interference patterns of water waves in a bucket implement the operations $\mathbf{A}\mathbf{x}(t)$ and $f(\cdot)$ in real time, while external sensors capture the instantaneous state $\mathbf{x}(t)$ (Fig. 1(b)) [9, 37]. By exploiting the medium’s native parallelism and energy efficiency, PRC offers a powerful neuromorphic alternative to digital computation.
A key innovation in PRC is the use of time-multiplexing to generate high-dimensional reservoir states with minimal hardware, often utilizing just a single nonlinear node (Fig. 1(c)) [8]. In this approach, during each clock period of length $T$, the input segment is multiplied by a masking sequence $\mathbf{m}$, and the element’s output is sampled at $M$ equidistant instants. The resulting vector
@@
constitutes an MM‑node virtual reservoir. The evolution of the reservoir state is governed by the interaction of the masked input and the reservoir’s inherent nonlinear dynamics and memory [39].
A common implementation of time-multiplexing involves an external feedback loop with delay $\tau$ around the nonlinear node (e.g., a semiconductor laser or an analogue circuit). Partitioning the delay into $M$ sub‑intervals of length $\Delta=\tau/M$ defines the $M$ virtual nodes [9, 40]. Sampling the output every $\Delta$ seconds constructs $\mathbf{x}(t)$, while the recirculating signal provides the necessary fading memory for tasks such as speech recognition and chaotic time‑series forecasting. Additional delay lines or delayed taps at the input or readout can further enrich the dynamics or extend the memory horizon without increasing the number of physical components [41].
@@ III Molecular Communication Channel as a Physical Reservoir
We consider a single-link MC system consisting of a point transmitter and a spherical receiver, as illustrated in Fig. 1(d). The transmitter releases a specific number of ligands (molecules) instantaneously at predefined intervals into a fluidic medium. These ligands propagate via free diffusion, and a fraction eventually reaches the receiver, where they bind reversibly to surface receptors under given reaction kinetics. We now detail how this canonical MC system is formalized as a physical reservoir computer, utilizing the time-varying fraction of bound receptors as the dynamical state.
Let the transmitter release $I(n)$ molecules at discrete symbol times $nT$, where $n\in\{0,1,2,\dots\}$, and $T$ is the symbol interval. We assume a 3D environment where the ligand concentration $c(\mathbf{x},t)$ evolves according to Fick’s second law of diffusion:
@@
where $D$ is the diffusion coefficient of ligands. We impose the far-field boundary condition $c(\mathbf{x},t)\to 0$ as $\|\mathbf{x}\|\to\infty$.
Assuming the receiver is fixed at a distance $d$ significantly larger than its radius $r_{\mathrm{rx}}$ ($d\gg r_{\mathrm{rx}}$), we can approximate the receiver as a point when calculating the channel impulse response. The impulse response at distance $d$ for a pulse released at $t=0$ is then given by the 3D diffusion kernel:
$h(t)\;=\;\frac{1}{(4\pi Dt)^{3/2}}\exp\!\Bigl(-\frac{d^{2}}{4Dt}\Bigr)\quad(t>0).$ (4)
The local ligand concentration at the receiver, $c_{\mathrm{R}}(t)$, resulting from the sequence of transmissions $I(n)$ is the superposition:
$c_{\mathrm{R}}(t)\;=\;\sum_{n=0}^{\infty}I(n)\,h\bigl(t-nT\bigr).$ (5)
The long tail of the diffusion kernel $h(t)$ inherently introduces ISI, providing the channel with signigicant memory. In numerical simulations, this tail is often truncated to a finite memory length [42].
Figure 2: (a) An instance from particle-based stochastic simulations of the MC channel with Smoldyn. Receptor occupation ratios for numerical and stochastic simulations of the MC system for (b) Mackey–Glass series input, (c) NARMA10 series input, (d) Mackey–Glass series input, post filtering (smoothing), (e) NARMA10 series input, post filtering.
The spherical receiver is equipped with $N_{R}$ identical, monovalent receptors. The interaction between ligands ($L$) and unbound receptors ($R$) is modeled as a reversible reaction forming a ligand-receptor complex ($LR$):
@@
At the mean-field level, assuming the local ligand concentration is not significantly depleted by binding, the time evolution of the mean fraction of bound receptors, $b(t)=n_{\mathrm{b}}(t)/N_{\mathrm{R}}$, can be approximated by
$\frac{d\,b(t)}{dt}\;=\;k_{\mathrm{on}}\,c_{\mathrm{R}}(t)\,\bigl[1-b(t)\bigr]\;-\;k_{\mathrm{off}}\,b(t),$ (6)
where $k_{\mathrm{on}}$ and $k_{\mathrm{off}}$ are the association (binding) and dissociation (unbinding) rate constants, respectively. This equation captures the essential nonlinearity (due to the saturation term [1b(t)1-b(t)]) and the memory (due to the kinetics) of the receiver response.
Equation (6) provides a deterministic, mean-field approximation of the receptor dynamics. This model is computationally efficient and allows for extensive parametric analysis. However, MC systems are inherently stochastic due to the discrete nature of molecules and the probabilistic nature of diffusion and binding events. To validate the mean-field model and assess the impact of molecular noise, we also employ particle-based spatial stochastic simulations using the Smoldyn simulator (Fig. 2) [43] in Section IV. Fig. 2(a) shows a snapshot from a Smoldyn simulation, illustrating the diffusion of ligands from the point source to the spherical receiver. Figs. 2(b) and 2(c) compare the evolution of the receptor occupation ratio $b(t)$, i.e., the reservoir’s internal state, for the deterministic model and the stochastic simulation when driven by two standard chaotic time series (Mackey–Glass and NARMA10). The stochastic simulations exhibit significant noise. However, when a moving-average filter ($2000$ ms window) is applied (Figs. 2(d) and 2(e)), it is evident that the deterministic model accurately captures the mean behavior of the stochastic system, justifying its use for analyzing the fundamental computational dynamics.
In the RC framework, the input time series $\{u(n)\}$ must be mapped into a high-dimensional, nonlinear dynamical state $\mathbf{x}_{n}$. We encode the input into the number of released molecules at time $nT$:
@@
where $\alpha$ is a scaling factor. The MC channel naturally imposes a complex, memory-rich transformation on this input sequence, resulting in the continuous-time scalar output $b(t)$.
To construct a high-dimensional reservoir state from $b(t)$, we employ time-multiplexing. Within each symbol interval [nT,(n+1)T][nT,\,(n+1)T], we sample $b(t)$ at $M$ distinct time offsets $\{\tau_{1},\tau_{2},\dots,\tau_{M}\}$, yielding the reservoir state vector:
@@
This process creates $M$ virtual nodes per symbol, akin to time-multiplexing used in delay-based RC, where a single nonlinear node is sampled at multiple, equally spaced intervals to emulate a high-dimensional network. In a typical delay-based RC, an external mask is required to diversify the dynamics of the virtual nodes, and an explicit delay loop provides memory. In the MC reservoir, these requirements are fulfilled intrinsically by the physical channel. The diffusion process inherently reshapes each molecular pulse in time, acting as a natural, dynamic mask. The resulting ISI from diffusion, combined with the memory from receptor kinetics, provides the necessary fading memory. Moreover, the receptor (receiver response) nonlinearity ensures that the temporal samples are dynamically rich and diverse.
The final step involves training a linear readout to map the reservoir state $\mathbf{x}_{n}$ to the desired output $y_{n}$ (e.g., a future value of the time series). The prediction $\hat{y}_{n}$ is given by:
@@
Given $N_{\mathrm{train}}$ training samples $\{(\mathbf{x}_{n},\,y_{n})\}$, the readout weights $\mathbf{W}_{\mathrm{out}}$ are optimized using ridge regression to minimize the regularized mean square error:
@@
where $\mathbf{X}$ is the state matrix (whose $n$-th row is $\mathbf{x}_{n}^{\top}$), $\mathbf{y}$ is the target vector, and $\lambda$ is a regularization parameter. The closed-form solution is given by the Moore–Penrose pseudoinverse:
@@
Importantly, only $\mathbf{W}_{\mathrm{out}}$ is trained, while the underlying physical dynamics of the MC channel remain fixed.